I wanna have a spreadsheet in excel (b & c).
5.19 Estimates have been presented to Holly Farms, which is considering two environmental chambers for a project that will detail laboratory confirma- tions of on-line bacteria tests in chicken meat for of E coli 0157:H7 and Listeriamono- cytogenes. (a) If the project will last for 6 years and 1-10% per year, perform an AW evaluation to determine which chamber is more economical (b) Chamber D103 can be purchased with differ- ent options and, therefore, at different installed costs. They range from $300,000 to $500,000. Will the selection change if one of these other models is installed? (c) Use single-cell spread- sheet functions to solve part (b). the presence Chamber D103 Chamber 490G Installed cost, S -400,000 -250,000 Annual operating cost $ per year Salvage value at 10% of P, $ -4000 -3000 40,000 25,000 Life, years
Answer #1

a) i= 10% and project life = 6 years then

The net equivalent annual worth (AW) of the Chamber D103 = -400,000-4000(A/F, 10%, 6) +40000(A/F, 10%,6)

                                                                 = -400000-4000* .1296 +40000* .1296 = $-395334.4

The net equivalent annual worth (AW) of the Chamber 490G= -250,000-3000(A/F, 10%, 6) +25000(A/F, 10%,6)

                                                                 = -250000-3000* .1296 +25000*.1296

                                                                = $--247148.8

As net AW of Chamber D103 is lower than net AW of Chamber 490G, therefore, Chamber 490G is economical.

b) Service life of Chamber D103= 3 years

The net equivalent annual worth (AW) of the Chamber D103 = -400,000-4000(A/F, 10%, 3) +40000(A/F, 10%,3)

                                                                 = -400000-4000* .3021 +40000* .3021 = $-389124.4

Service life of Chamber 490G= 2 years

The net equivalent annual worth (AW) of the Chamber 490G= -250,000-3000(A/F, 10%, 2) +25000(A/F, 10%,2)

                                                                 = -250000-3000* .4762 +25000* .4762 = $-239523.6

The net AW of Chamber 490G is higher than net AW of Chamber D103. Therefore, the selection as shown in part (a) will not change in this case also.

c) Excel Spreadsheet depiction of net annual worth of Chamber D103 using annuity table for discrete compounding:

Chamber D103
Annual Worth
Installed cost -400000 -400000
Annual operating cost -4000 -1208.4
Salvage value 40000 12084
total -389124.4

Excel Spreadsheet depiction of net annual worth of Chamber 490G using annuity table for discrete compounding:

Chamber 490G
Annual Worth
Installed cost -250000 -250000
Annual operating cost -3000 -1428.6
Salvage value 25000 11905
total -239523.6